a drug is known to be ineffective after it has decomposed 25% the original concentration of a sample was 500 unkits/ml when analyzed 20 months later the concentration was found to be 450 units /ml assuming that decomposition is of 1st order what will be the expiry time of the drug?
 

Grade:11

1 Answers

BIKRAM
6 years ago
ln(A0/A)= Kt : K=(1/20×ln(500/450)=5.268×10-3 mon-1
A= 500-500×(25/100)=375
expiry date (t)= (1/K)ln(500/375)= 54.6 month